# The discrete periodic Pitman transform: invariances, braid relations, and Burke properties

### v1

@@ Abstract.
We develop the theory of the discrete periodic Pitman transform, first introduced by Corwin, Gu, and the fifth author. We prove that, for polymers in a periodic environment, single-path and multi-path partition functions are preserved under the action of this transform on the weights in the polymer model. As a corollary, we prove that the discrete periodic Pitman transform satisfies the same braid relations that are satisfied for the full-line Pitman transform, shown by Biane, Bougerol, and O’Connell. Combined with a new inhomogeneous Burke property for the periodic Pitman transform, we prove a multi-path invariance result for the periodic inverse-gamma polymer under permutations of the column parameters. In the limit to the full-line case, we obtain a multi-path extension of a recent invariance result of Bates, Emrah, Martin, Seppäläinen, and the fifth author, in both positive and zero-temperature. Additionally, we give a combinatorial description of the distribution of the $2$-component jointly invariant measures for a discrete-time Markov chain.
@@ 1. Introduction
The purpose of this paper is to demonstrate the rich algebraic and probabilistic structure of a periodic version of the Pitman transform, recently introduced in [CGS24]. Our first main result, Theorem 1.4, gives an invariance result for directed polymers in a periodic environment under a periodic Pitman transformation of the column weights. Previous results of this flavor in a non-periodic setting have appeared in several contexts and are discussed after the statement of the theorem. The key machinery is the encoding of geometric RSK in terms of matrix relations from [NY04]. Corwin [Cor21] previously used these tools to prove the analogous invariance for full-space polymers. The next main result, Theorem 1.6, uses this polymer invariance result to prove braid relations for the periodic Pitman transform, analogous to the relations enjoyed by the full-line Pitman transform [BBO05]. More specifically, we consider a collection of operators on the space $(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ of bi-infinite sequences of $N$-periodic vectors and operators $\mathcal{P}_{k}$, which replace the $k$th and k+1k+1st vector with their Pitman transform. The braid relations, along with the fact that each $\mathcal{P}_{k}$ is an involution (proved in [CGS24], see Lemma A.2) shows that this collection of operators generates a copy of the infinite symmetric group. Both of these main results pass to zero-temperature limits, stated in Theorems 1.8 and 1.9. We then utilize the invariance of partition functions and a new inhomogeneous Burke property (Propositions 5.1, 5.2, and 5.3) to prove a distributional invariance of the inverse-gamma polymer partition function under the permutation of the row and column parameters in both the periodic and full-line cases (Theorems 1.10 and 1.11). In the full-line case, this gives a multi-path and positive temperature extension of a recent result of Bates, Emrah, Martin, Seppäläinen, and the fifth author [BEM+25b]. This result bears similarity to an invariance found previously by Dauvergne [Dau22]. See the discussion before and after Theorem 1.11 for more on the connection to other works.
Beginning with the work of Ferrari and Martin in particle systems [FM07, FM06, FM09], and adapted to last-passage percolation, polymers, and their scaling limits in [FS20, SS23, Bus24, BSS24, BFS25, GRASS25, CGS24], Pitman transforms have been used to describe the distribution of jointly invariant or multi-type invariant measures in these models. Indeed, the work [CGS24] introduced the periodic Pitman transform to describe the jointly invariant measures for the KPZ equation with periodic noise. A similar object was previously used in [FM07] to find the stationary measure for multi-species TASEP on a ring (see [CGS24, Appendix F] for more of a discussion on this connection). In Theorem 1.12, we develop a similar description of the jointly invariant measures for a discrete-time Markov chain in zero-temperature. Then, in Theorem 1.13, we give a combinatorial description of this invariant measure. We can obtain these joint stationary measures in the full-line settings by taking appropriate scaling limits of the measure we study in Theorem 1.13, and we hope that our description will lead to further understanding in future work.
@@ 1.1. The positive-temperature discrete periodic Pitman transform
For $N\in\mathbb{N}$, let $\mathbb{R}^{\mathbb{Z}_{N}}$ denote the set of $N$-tuples, together with periodic indexing. That is, for $\mathbf{X}\in\mathbb{R}^{\mathbb{Z}_{N}}$, $\mathbf{X}=(X_{1},X_{2},\ldots,X_{N})$, where the indices represent elements of $\mathbb{Z}_{N}$ (see Figure 1). For a vector $\mathbf{X}\in\mathbb{R}^{\mathbb{Z}_{N}}$ and indices $i,j\in\mathbb{Z}_{N}$, we set
@@
We define the positive-temperature discrete periodic Pitman transform as the map $\mathcal{P}:\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}\to\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}$ defined by
@@ 1.2. Polymers on $\mathbb{Z}^{2}$ and periodic environments
$Z^{\beta,\mathbf{W}}(n,m\mid i,j):=\sum_{\pi\in\Pi_{(i,j),(n,m)}}\prod_{(r,\ell)\in\pi}e^{\beta W_{(r,\ell)}}$ (1.7)
@@
$Z^{\beta,\mathbf{W}}(V\mid U):=\sum_{\pi\in\Pi_{U,V}}\prod_{(r,\ell)\in\pi}e^{\beta W_{(r,\ell)}}$ (1.8)
@@ 1.3. Invariance of periodic polymer partition functions
Let $N\in\mathbb{N}$, and let $\mathbf{W}=(\mathbf{W}_{k}\in\mathbb{R}^{\mathbb{Z}_{N}}\mid k\in\mathbb{Z})=(W_{k,\ell}\mid k\in\mathbb{Z},\ell\in\mathbb{Z}_{N})$ be a collection of real-valued weights. Extend these periodically to weights on $\mathbb{Z}\times\mathbb{Z}$ by the condition $W_{k,\ell}=W_{j,\ell}$ whenever $k\equiv j\mod N$. If $k_{1},k_{2},\ldots,k_{m}\in\mathbb{Z}$ and the sets of initial points $U=\bigl{(}\bigl{(}a_{i},b_{i})_{i\in\llbracket 1,m\rrbracket}\bigr{)}$ and end points $V=\bigl{(}(c_{i},d_{i})_{i\in\llbracket 1,m\rrbracket}\bigr{)}$ satisfy $(U,V)\in\Psi$ and, for all $i,j\in\llbracket 1,m\rrbracket$, $a_{i}\neq k_{j}+1$ and $c_{i}\neq k_{j}$, then,
@@
Figure 3. A depiction of the polymer invariance in Theorem 1.4. Each column is associated to a periodic vector of weights, $\mathbf{W}_{k}$. The horizontal lines are used to indicate a periodic environment for the polymer with $N=4$. In this example, we take $m=1$ and consider the action of the operator $\mathcal{P}_{3}$, which replaces the weights in columns $3$ and $4$ with $T(\mathbf{W}_{4},\mathbf{W}_{3})$ and $D(\mathbf{W}_{4},\mathbf{W}_{3})$, respectively. The partition function for each pair of points and the multi-path partition function for this collection of $3$ pairs are both preserved under this transformation of the weights.
@@ 1.4. Braid relations
Theorem 1.6 is proved in Section 3. The key is the matrix relation in Corollary 2.5 below, which is also used to prove Theorem 1.4. When written out, the braid relation consists of three equalities ((3.1)-(3.3) below). The third of these equalities was previously proved in [CGS24, Proposition 5.13]. With that as our starting point, we iterate the matrix relation in Corollary 2.5 in two different ways to obtain and use the interpretation of the matrix product as the polymer partition function to obtain the other two equalities.
@@ 1.5. Zero-temperature limits
One immediately sees that we can obtain $\widetilde{\mathcal{P}}$ as a limiting case of the positive-temperature discrete periodic Pitman transform:
@@
For real-valued weights $\mathbf{W}=\bigl{(}W_{(i,j)}\in\mathbb{R}\mid i,j\in\mathbb{Z}\bigr{)}$ corresponding to the vertices of the directed lattice $\mathbb{Z}^{2}$, we define the last-passage percolation (LPP) time between the points $(i,j)$ and $(n,m)$ in $\mathbb{Z}^{2}$ as the maximal sum over weights along an up-right path between $(i,j)$ and $(n,m)$. That is,
$G^{\mathbf{W}}(n,m\mid i,j)\coloneqq\max\limits_{\pi\in\Pi_{(i,j),(n,m)}}\sum_{(r,\ell)\in\pi}W_{(r,\ell)}$
We similarly define multi-path last-passage times analogously as in (1.8):
@@
With this notation defined, we get the following zero-temperature analogues of Theorems 1.4 and 1.6:
@@
The same invariance in Theorem 1.4 holds, $Z^{\mathbf{W}}$ is replaced by $G^{\mathbf{W}}$, and $T,D$ are replaced by $\widetilde{T},\widetilde{D}$.
@@
For $k\in\mathbb{Z}$, define the map $\widetilde{\mathcal{P}}_{k}:(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}\to(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ by (1.9), with $T,D$ replaced by $\widetilde{T},\widetilde{D}$. Then, the same braid relations in Theorem 1.6 hold.
The details of the limit transitions to Theorems 1.8 and 1.9 are handled in Section 4.
1.6. Permutation invariance
In this section, we consider an inhomogeneous, periodic inverse-gamma polymer. Let $\lambda>0$ and $\mathbf{a}\in\mathbb{R}^{\mathbb{Z}}$ and $\mathbf{b}\in\mathbb{R}^{\mathbb{Z}_{N}}$ be sequences of real numbers such that $a_{i}+b_{j}>0$ for $i\in\mathbb{Z}$ and $j\in\mathbb{Z}_{N}$. We take independent weights $\mathbf{W}^{\mathbf{a},\mathbf{b}}=\bigl{(}W_{(i,j)}\mid i\in\mathbb{Z},\ j\in\mathbb{Z}_{N}\bigr{)}$ such that $W_{(i,j)}\sim\text{Log-Inv-Gamma}(a_{i}+b_{j},\lambda)$ for $i\in\mathbb{Z}$ and $j\in\mathbb{Z}_{N}$ (see the notation section below). Extend these periodically to weights on $\mathbb{Z}\times\mathbb{Z}$. We call a permutation $\sigma:\mathbb{Z}\to\mathbb{Z}$ a finite permutation if $\sigma$ fixes all but finitely many coordinates. For a finite permutation $\sigma$, we let $\mathcal{U}^{\sigma}$ be the set of tuples $(U,V)\in\Psi$, such that
[…]

### v2

@@ Abstract.
We develop the theory of the discrete periodic Pitman transform, first introduced by Corwin, Gu, and the fifth author. We prove that, for polymers in a periodic environment, single-path and multi-path partition functions are preserved under the action of this transform on the weights in the polymer model. As a corollary, we prove that the discrete periodic Pitman transform satisfies the same braid relations that are satisfied for the full-line Pitman transform shown by Biane, Bougerol, and O’Connell. Combined with a new inhomogeneous Burke property for the periodic Pitman transform, we prove a multi-path invariance result for the periodic inverse-gamma polymer under permutations of the column parameters. In the limit to the full-line case, we obtain a multi-path extension of a recent invariance result of Bates, Emrah, Martin, Seppäläinen, and the fifth author, in both positive and zero-temperature.
@@ 1. Introduction
The purpose of this paper is to demonstrate the rich algebraic and probabilistic structure of a periodic version of the Pitman transform, recently introduced in [CGS24]. Our first main result, Theorem 1.4, gives an invariance result for directed polymers in a periodic environment under a periodic Pitman transformation of the column weights. Previous results of this flavor in a non-periodic setting have appeared in several contexts and are discussed after the statement of the theorem. The key machinery is the encoding of geometric RSK in terms of matrix relations from [NY04]. Corwin [Cor21] previously used these tools to prove the analogous invariance for full-space polymers. The next main result, Theorem 1.6, uses this polymer invariance result to prove braid relations for the periodic Pitman transform, analogous to the relations enjoyed by the full-line Pitman transform [BBO05]. More specifically, we consider a collection of operators on the space $(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ of bi-infinite sequences of $N$-periodic vectors and operators $\mathcal{P}_{k}$, which replace the $k$th and k+1k+1st vector with their Pitman transform. The braid relations, along with the fact that each $\mathcal{P}_{k}$ is an involution (proved in [CGS24], see Lemma A.2) shows that this collection of operators generates a copy of the infinite symmetric group. We then utilize the invariance of partition functions and a new inhomogeneous Burke property (Proposition 4.1) to prove a distributional invariance of the inverse-gamma polymer partition function under the permutation of the row and column parameters in both the periodic and full-line cases (Theorems 1.7 and 1.8). In the full-line case, this gives a multi-path and positive temperature extension of a recent result of Bates, Emrah, Martin, Seppäläinen, and the fifth author [BEM+25b]. This result bears similarity to an invariance found previously by Dauvergne [Dau22]. See the discussion before and after Theorem 1.8 for more on the connection to other works. Each of our main results has an analogue in zero-temperature. These results are stated in Section 1.6.
@@ 1.1. The positive-temperature discrete periodic Pitman transform
For $N\in\mathbb{N}$, let $\mathbb{R}^{\mathbb{Z}_{N}}$ denote the set of $N$-tuples, together with periodic indexing. That is, for $\mathbf{X}\in\mathbb{R}^{\mathbb{Z}_{N}}$, $\mathbf{X}=(X_{1},X_{2},\ldots,X_{N})$, where the indices represent elements of $\mathbb{Z}_{N}$ with addition modulo $N$ (see Figure 1). For a vector $\mathbf{X}\in\mathbb{R}^{\mathbb{Z}_{N}}$ and indices $i,j\in\mathbb{Z}_{N}$, we set
@@
The positive-temperature discrete periodic Pitman transform is the map $\mathcal{P}:\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}\to\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}$ defined by
@@ 1.2. Polymers on $\mathbb{Z}^{2}$ and periodic environments
$Z^{\beta,\mathbf{W}}(n,m\mid i,j):=\sum_{\pi\in\Pi_{(i,j),(n,m)}}\prod_{(r,\ell)\in\pi}e^{\beta W_{(r,\ell)}}.$ (1.7)
@@
$Z^{\beta,\mathbf{W}}(V\mid U):=\sum_{\pi\in\Pi_{U,V}}\prod_{(r,\ell)\in\pi}e^{\beta W_{(r,\ell)}}.$ (1.8)
@@ 1.3. Invariance of periodic polymer partition functions
Let $N\in\mathbb{N}$, and let $\mathbf{W}=(\mathbf{W}_{k}\in\mathbb{R}^{\mathbb{Z}_{N}}\mid k\in\mathbb{Z})=(W_{k,\ell}\mid k\in\mathbb{Z},\ell\in\mathbb{Z}_{N})$ be a collection of real-valued weights. Extend these periodically to weights on $\mathbb{Z}\times\mathbb{Z}$ by the condition $W_{k,j}=W_{k,\ell}$ whenever $j\equiv\ell\mod N$. If $k_{1},k_{2},\ldots,k_{m}\in\mathbb{Z}$ and the sets of initial points $U=\bigl{(}\bigl{(}a_{i},b_{i})_{i\in\llbracket 1,m\rrbracket}\bigr{)}$ and end points $V=\bigl{(}(c_{i},d_{i})_{i\in\llbracket 1,m\rrbracket}\bigr{)}$ satisfy $(U,V)\in\Psi$ and, for all $i,j\in\llbracket 1,m\rrbracket$, $a_{i}\neq k_{j}+1$ and $c_{i}\neq k_{j}$, then,
@@
Figure 3. A depiction of the polymer invariance in Theorem 1.4. Each column is associated to a periodic vector of weights, $\mathbf{W}_{k}$. The horizontal lines are used to indicate a periodic environment for the polymer with $N=4$. In this example, we take $m=1$ and consider the action of the operator $\mathcal{P}_{3}$, which replaces the weights in columns $3$ and $4$ with $T(\mathbf{W}_{4},\mathbf{W}_{3})$ and $D(\mathbf{W}_{4},\mathbf{W}_{3})$, respectively. The partition function for each pair of points and the multi-path partition function for this collection of 33 pairs are both preserved under this transformation of the weights.
@@ 1.4. Braid relations
Theorem 1.6 is proved in Section 3. The key is the matrix relation in Corollary 2.5 below, which is also used to prove Theorem 1.4. When written out, the braid relation consists of three equalities ((3.1)-(3.3) below). The third of these equalities was previously proved in [CGS24, Proposition 5.13]. With that as our starting point, we iterate the matrix relation in Corollary 2.5 in two different ways to obtain and use the interpretation of the matrix product as the polymer partition function to obtain the other two equalities. Specifically, Lemma 3.1 is the key innovation that allows us to complete the proof.
1.5. Permutation invariance
In this section, we consider an inhomogeneous, periodic inverse-gamma polymer. Let $\lambda>0$ and $\mathbf{a}\in\mathbb{R}^{\mathbb{Z}}$ and $\mathbf{b}\in\mathbb{R}^{\mathbb{Z}_{N}}$ be sequences of real numbers such that $a_{i}+b_{j}>0$ for $i\in\mathbb{Z}$ and $j\in\mathbb{Z}_{N}$. We take independent weights $\mathbf{W}^{\mathbf{a},\mathbf{b}}=\bigl{(}W_{(i,j)}\mid i\in\mathbb{Z},\ j\in\mathbb{Z}_{N}\bigr{)}$ such that $W_{(i,j)}\sim\text{Log-Inv-Gamma}(a_{i}+b_{j},\lambda)$ for $i\in\mathbb{Z}$ and $j\in\mathbb{Z}_{N}$ (see Definition 1.15 below). Extend these periodically to weights on $\mathbb{Z}\times\mathbb{Z}$. We call a permutation $\sigma:\mathbb{Z}\to\mathbb{Z}$ a finite permutation if $\sigma$ fixes all but finitely many coordinates. For a finite permutation $\sigma$, we let $\mathcal{U}^{\sigma}$ be the set of tuples $(U,V)\in\Psi$, such that
[…]
Theorem 1.7.
where $\sigma\mathbf{a}=(a_{\sigma(i)})_{i\in\mathbb{Z}}$.
[…]
The following is an extension of the recent result [BEM+25a, Theorem 3.1] to multi-paths and to the positive-temperature case. A sketch of the proof for the extension to multi-paths in zero temperature was previously discussed in [BEM+25a, Remark 4.3]. The technique follows the same outline as developed in [BEM+25a], as we use a coupling of the weights given by applying a Pitman transform to neighboring rows and columns. However, the coupling we use in this proof is different because we treat the weights as periodic outside a large box and apply the periodic Pitman transform. The algebra involved in this positive-temperature case is also different.
Theorem 1.8.
where $\sigma\mathbf{a}=(a_{\sigma(i)})_{i\in\mathbb{Z}}$ and $\tau\mathbf{b}=(b_{\tau(j)})_{j\in\mathbb{Z}}$.
Theorems 1.7 and 1.8 are proved in Section 4. When the initial point is fixed, invariances of this flavor can be deduced from the symmetry of Schur functions and Whittaker functions (see, for example, [BP08, COSZ14, OSZ14, BC14]). A similar type of invariance for last-passage percolation was discovered in the work of Dauvergne [Dau22, Theorem 1.5], motivated by a shift-invariance result previously proved in Borodin, Gorin, and Wheeler [BGW22]. Connections between these results are discussed further in [BEM+25a, Section 3].
@@ 1.6. Zero-temperature limits
The mapping $\widetilde{D}$ is similar to, but not equivalent to a mapping used to describe the invariant measures for multi-type TASEP by Ferrari and Martin [FM07]. See [CGS24, Appendix F] for more on this connection. One immediately sees that we can obtain $\widetilde{\mathcal{P}}$ as a limiting case of the positive-temperature discrete periodic Pitman transform:
@@
For real-valued weights $\mathbf{W}=\bigl{(}W_{(i,j)}\in\mathbb{R}\mid i,j\in\mathbb{Z}\bigr{)}$ corresponding to the vertices of the directed lattice $\mathbb{Z}^{2}$, we define the last-passage percolation (LPP) time for both a single pair of points and multi-paths analogously as in (1.7)-(1.8):
@@
For $k\in\mathbb{Z}$, define the map $\widetilde{\mathcal{P}}_{k}:(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}\to(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ by (1.9), with $T,D$ replaced by $\widetilde{T},\widetilde{D}$.
With this notation defined, we get the following zero-temperature analogues of our main theorems.
@@
The same invariance in Theorem 1.4 holds when $Z^{\mathbf{W}}$ is replaced by $G^{\mathbf{W}}$, and the maps $\mathcal{P}_{k_{j}}$ are replaced by $\widetilde{\mathcal{P}}_{k_{j}}$.
@@
The same braid relations in Theorem 1.6 hold for the maps $\widetilde{\mathcal{P}}_{k}$.